# RLHFuse: Efficient RLHF Training for Large Language Models with Inter- and Intra-Stage Fusion

### v1

@@ Abstract
Reinforcement Learning from Human Feedback (RLHF) stands as a pivotal post-training technique to enhance the alignment between LLMs and human preference. The workflow of RLHF typically involves several models and tasks in a series of distinct stages. Existing RLHF training systems view each task as the smallest execution unit and thus overlooking the opportunities for subtask-level optimizations. Due to the intrinsic nature of RLHF training, i.e., the data skewness in the generation stage, and the pipeline bubbles in the training stage, existing RLHF systems suffer from low GPU utilization in production deployments.
@@ 1 Introduction
Figure 1: Top: RLHF training workflow. Bottom: RLHF training iteration breakdown of existing RLHF Training frameworks.
@@
Unlike pre-training, RLHF training is significantly more complex in terms of system design, since it involves four different LLMs and three stages. As shown in the upper of Figure 1, each iteration begins with the primary LLM, referred to as the Actor model, which generates responses to a batch of input prompts sampled from the dataset (known as the generation stage). These generated responses are then assessed by three additional LLMs: the Reference model, the Critic model, and the Reward model, during the inference stage. Finally, the Actor and Critic models use the evaluation results to compute loss and update their parameters, completing the training stage. In summary, there are six tasks (nodes) in the training workflow. The data and weight dependencies (edges between nodes) create synchronization boundaries between tasks.
To enable efficient training of RLHF, existing frameworks have proposed a series of optimizations, such as selecting tailored parallel strategies for each task [13] and optimizing the data exchange during stage transitions [14]. However, these techniques fundamentally treat the RLHF workflow as a simple composition of individual tasks, failing to delve into the inherent characteristics and structure inside the tasks, thereby missing significant optimization opportunities. In production deployments, we have observed two issues within current frameworks that severely impact training efficiency, as indicated by the grey area at the bottom of Figure 1.
@@
To this end, we propose RLHFuse, the first work to employ subtask-level optimizations for RLHF training. By breaking down the task into finer-grained subtasks, RLHFuse opens up a new design space to fuse the stage execution. The core of RLHFuse is data-aware inter-stage fusion and model-aware intra-stage fusion. These two techniques leverage the unique system characteristics of RLHF training to address data skewness and pipeline bubbles, respectively.
@@ 2.1 Background
Training stage. All the samples are divided into several mini-batches. For each mini-batch, the Actor and Critic models first perform a forward pass, then calculate the loss using the results from the inference stage and perform a backward pass to update the parameters. The parameter update must occur before the forward pass of subsequent mini-batch, distinguishing this approach from gradient accumulation, which performs a single parameter update across mini-batches. After the Actor completes training on all mini-batches, it uses the up-to-date parameters for the generation in the next iteration.
@@
Data Parallelism (DP). DP divides training data along the batch dimension and sends each partition to a model replica. At the end of each iteration, all DP groups perform all-reduce communication on the gradients before updating parameters.
@@
Since the above parallelism methods are orthogonal, each with its own advantages, all three parallel strategies are employed together to scale training in practice [16]. We denote the 3D-parallelism strategy by an integer triple $(dp,pp,tp)$, representing the degrees of DP, PP, and TP, respectively.
@@ 3 RLHFuse Overview
To this end, we present RLHFuse to address the above problems. With specific algorithm, model, and cluster configurations as input, RLHFuse first generates efficient parallel and deployment strategies for each task within the RLHF workflow. During runtime, RLHFuse addresses data skewness in the generation stage through data-aware inter-stage fusion (§4) and mitigates pipeline bubbles in the training stage with model-aware intra-stage fusion (§5). Additionally, RLHFuse adopts a series of system optimizations (§6) tailored for RLHF training. Here we provide a brief overview of RLHFuse as shown in Figure 4.
Execution plan generator. We formalize the training iteration of RLHF workflow as a dataflow graph $\mathcal{G}(\mathcal{V},\mathcal{E})$ following previous work [13]. A node $v_{i}\in\mathcal{V}$ represents the $i$-th task in the training iteration. An edge $(v,v^{\prime})\in\mathcal{E}$ indicates the data or weights dependency. For each $v_{i}$, RLHFuse will assign it a tailored parallel strategy $P_{i}=(dp,pp,tp)$ based on the model and cluster configurations to maximize the GPU utilization. To deploy the task onto the physical cluster, we introduce the concept of device mesh, which is the unit for executing an individual task. A device mesh $D=(N,M)$ is defined as a two-dimensional grid of GPUs, covering $N$ nodes where each is equipped with $M$ devices. RLHFuse will designate a device mesh $D_{i}$ to each task $v_{i}$ which can be executed with parallel strategy $P_{i}$. Finally, the execution plan generator will output an augmented dataflow graph $\mathcal{G}^{*}(\mathcal{V},\mathcal{E},\mathcal{P},\mathcal{D})$ which can be deployed on the physical cluster.
@@
Runtime executor. Given the augmented dataflow graph $\mathcal{G}^{*}(\mathcal{V},\mathcal{E},\mathcal{P},\mathcal{D})$, the runtime executor will launch each task with its corresponding parallel strategy on the designated device mesh following the workflow dependency. During the generation stage, there are multiple model instances, each managing complete model weights and part of the samples. RLHFuse actively monitors each generation instance and triggers migration when the number of remaining samples is below a particular threshold. The remaining samples will be migrated to some dedicated generation instances which are designated to handle long-tailed samples. Subsequently, the resources of the original generation instances are repurposed to launch inference tasks, thereby allowing for the overlap of inference tasks with the generation of long-tailed samples. In the training stage, it will fuse the two training tasks with a bi-directional pipeline schedule to mitigate the pipeline bubbles and optimize the activation memory usage. Throughout the iteration, the runtime executor will handle the weight redistribution and data transmission between different tasks to ensure synchronous training semantics.
@@ 4.1 Opportunities and Challenges
Challenges. With the inter-stage fusion, the problem now lies in generating an efficient fused execution plan that decides the migration timing, destination, and mechanism based on the augmented dataflow graph $\mathcal{G}^{*}$. These three factors significantly affect the overall execution time of the two stages, and therefore must be chosen carefully.
@@ 6 System Implementation and Optimizations
Execution plan generator. Optimizing the parallel strategy for LLM training is a well-studied problem [33, 16, 34, 13]. Due to the deterministic nature of LLM computation, the execution time and memory cost can be accurately modeled through minimal profiling. An optimal solution can then be found through optimizations [33, 25, 34]. The generation and inference task involves only the forward pass, which is a subproblem of the LLM training. RLHFuse adopts a similar model-then-optimize approach. We build a simulator to accurately estimate the runtime statistics of the tasks under a specific parallel strategy and workload. Then we follow the guidelines in the Megatron-LM [16] paper to prune the design space and brute-force search the optimal strategy.
Runtime Executor. For inter-stage fusion, we use the simulator mentioned above to choose the best migration threshold $R_{t}$ to minimize the fused execution time under a specific workload pattern. For intra-stage fusion, we use MPI [35] to parallelize the simulated annealing with different random seeds. After finding the optimal solution, it will generate the sequences of NCCL [36] operations on each device following the schedule. The runtime will follow the instruction flow to execute the fused pipeline schedule.
System optimizations. Additionally, we adopt a series of system optimizations tailored for RLHF training.
@@
Task switching. The primary overhead during task switching lies in model weights reloading. For the Actor and Critic models, the latest weights after the training stage need to be redistributed to the appropriate devices according to the new parallel strategy for generation or inference. This communication pattern is well-studied [40] and we adopt a similar all-to-all approach to minimize cross-node communication. For the Ref and RW models whose weights remain unchanged, we keep them in CPU memory and swap them into GPU memory as needed. The swapping can overlap with the computation of previous tasks. With these optimizations, the task switching overhead is negligible compared to the total iteration time, as shown in §7.2.
@@ 7.1 End-to-end Experiment
7.2 Breakdown Analysis
To delve deeper into the performance improvement of RLHFuse, we show its RLHF iteration breakdown and compare it with RLHFuse-Base. As shown in Figure 8, each row represents a generation setting, and each column represents a model setting X/Y. We divide one RLHF iteration into three main parts: the generation plus inference stage (Gen.+Inf.), the training stage (Train), and all other overheads (Others), such as the data transmission and weights redistribution time.
@@ 7.2 Breakdown Analysis
7.3 Ablation Study
In this section, we conduct ablation study to show the effectiveness of RLHFuse’s key techniques in §4 and §5.
@@ 7.3 Ablation Study
7.4 Case Study
In this section, we showcase a fused pipeline schedule produced by RLHFuse for the 65B/33B model setting. The upper of Figure 10 shows the GPU execution timeline, with forward and backward tasks in light and dark colors, respectively. With intra-stage fusion, RLHFuse strategically fuses one 65B model of 16 PP stages (blue grids) with two 33B models that each has 8 PP stages (pink grids) to fill the pipeline bubbles of each other in reversed pipeline directions. Ultimately, the execution time of this fused pipeline schedule is the same as individually executing the 65B model using 1F1B schedule, which means we completely overlap the training of the 33B model, achieving the theoretical lower bound. Meanwhile, as shown in the bottom of Figure 10, the peak activation memory usage also achieves the lower bound of executing the two models in serial with 1F1B schedule. This complex schedule, which achieves optimal performance in both time and memory usage, is difficult to design manually, further demonstrating the effectiveness of our fused pipeline schedule generation algorithm proposed in §5.2.

### v2

@@ Abstract
Reinforcement Learning from Human Feedback (RLHF) enhances the alignment between LLMs and human preference. The workflow of RLHF typically involves several models and tasks in a series of distinct stages. Existing RLHF training systems view each task as the smallest execution unit and thus overlooking the opportunities for subtask-level optimizations. Due to the intrinsic nature of RLHF training, i.e., the data skewness in the generation stage, and the pipeline bubbles in the training stage, existing RLHF systems suffer from low GPU utilization in production deployments.
@@ 1 Introduction
Figure 1: RLHF iteration breakdown of existing frameworks.
@@
Unlike pre-training, RLHF training is significantly more complex in terms of system design, since it involves four different LLMs, three stages, and six tasks in total. As shown in Figure 1, each iteration begins with the Actor model, which generates responses to a batch of input prompts (known as the generation stage). These responses are then passed to the Reference, Critic, and Reward model for processing, during the inference stage. Finally, the Actor and Critic models compute loss and update their parameters, completing the training stage. The data and weight dependencies create synchronization boundaries between stages.
To enable efficient training of RLHF, existing frameworks have proposed a series of optimizations, such as selecting tailored parallel strategies for each task [13] and optimizing the data exchange during stage transitions [14]. However, these techniques fundamentally treat the RLHF workflow as a simple composition of individual tasks, failing to delve into the inherent characteristics and structure inside the tasks, thereby missing significant optimization opportunities. In production deployments, we have observed two issues within current frameworks that severely impact training efficiency, as indicated by the grey area in Figure 1.
@@
To this end, we propose RLHFuse, the first work to employ subtask-level optimizations for RLHF training. By breaking down the task into finer-grained subtasks, RLHFuse opens up a new design space to fuse the stage execution. This is orthogonal to the state-of-the-art RLHF framework ReaLHF [13] that optimizes the parallelism strategy for each task. RLHFuse exploits the optimizations opportunities at the subtask level with two new techniques, i.e., data-aware inter-stage fusion and model-aware intra-stage fusion. These two techniques leverage the unique system characteristics of RLHF training to address data skewness and pipeline bubbles, respectively.
@@ 2.1 Background
Training stage. In the PPO algorithm, unlike traditional LLM training, all the samples are divided into several mini-batches and the model weights need to update after training on each mini-batch. Concretely, for each mini-batch, the Actor and Critic models first perform a forward pass, then calculate the loss using the results from the inference stage and perform a backward pass to update the parameters. After the Actor completes training on all mini-batches, it uses the up-to-date parameters for the generation in the next iteration.
@@
Data Parallelism (DP). DP replicates the model weights and distribute the data among the replicas to execute in parallel.At the end of each iteration, all DP groups perform all-reduce communication on the gradients before updating parameters.
@@
Since the above parallelism methods have their own advantages, all three parallel strategies are used together to scale training in practice, such as Megatron-LM [16] and MegaScale [17] in pre-training [16] and ReaLHF [13] in RLHF training. We use an integer triple $(dp,pp,tp)$ to denote the degrees of DP, PP, and TP in 3D-parallelism.
@@ 3 RLHFuse Overview
To this end, we present RLHFuse to address data skewness in the generation stage through data-aware inter-stage fusion (§4) and mitigates pipeline bubbles in the training stage with model-aware intra-stage fusion (§5). Additionally, RLHFuse adopts a series of system optimizations (§6) tailored for RLHF training. Here we provide a brief overview of RLHFuse as shown in Figure 4.
Workflow. With specific algorithm, model, and cluster configurations as input, RLHFuse first configures efficient parallel and deployment strategies following the approach in ReaLHF [13]. For each task within the RLHF workflow, RLHFuse will assign it a tailored parallel strategy $P_{i}=(dp,pp,tp)$ to maximize the GPU utilization. To deploy the task onto the physical cluster, we introduce the concept of device mesh, which is the unit for executing an individual task. A device mesh $D=(N,M)$ is defined as a two-dimensional grid of GPUs, covering $N$ nodes where each is equipped with $M$ devices. RLHFuse will designate a device mesh $D_{i}$ to each task which can be executed with parallel strategy $P_{i}$. At runtime, RLHFuse will launch each task with its corresponding parallel strategy on the designated device mesh following the workflow dependency and handle the weight redistribution and data transmission between different tasks to ensure synchronous training semantics. Meanwhile, RLHFuse utilizes data-aware inter-stage fusion to fuse the generation and inference tasks and model-aware intra-stage fusion to fuse the training tasks.
Inter-stage fusion. As depicted in the left part of Figure 4, during the generation stage, there are multiple model instances, each managing complete model weights and part of the samples. RLHFuse actively monitors each generation instance and triggers migration when the number of remaining samples is below a particular threshold. The remaining samples will be migrated to some dedicated generation instances which are designated to handle long-tailed samples. Subsequently, the resources of the original generation instances are repurposed to launch inference tasks, thereby allowing for the overlap of inference tasks with the generation of long-tailed samples and fusing the execution of the two stages.
Intra-stage fusion. As shown in the right part of Figure 4, in the training stage, RLHFuse leverages the insight that pipeline execution can be complemented by another pipeline and generates a fused pipeline schedule based on the model size and parallel strategies of Actor and Critic model. It utilizes a light-weight schedule generation algorithm to minimize the pipeline bubbles and activation memory usage, thus improving the training efficiency.
@@ 4.1 Opportunities and Challenges
Challenges. With the inter-stage fusion, the problem now lies in generating an efficient fused execution plan that decides the migration timing, destination, and mechanism. These three factors significantly affect the overall execution time of the two stages, and therefore must be chosen carefully.
@@ 6 System Implementation and Optimizations
Parallelism strategy configuration. Optimizing the parallel strategy for LLM training is a well-studied problem [33, 16, 34, 13]. Due to the deterministic nature of LLM computation, the execution time and memory cost can be accurately modeled through minimal profiling. An optimal solution can then be found through optimizations [33, 25, 34, 13]. The generation and inference task involves only the forward pass, which is a subproblem of the LLM training. ReaLHF [13] is a state-of-the-art solution that optimizes the parallel strategy for RLHF training and RLHFuse adopts a similar model-then-optimize approach. We build a simulator to accurately estimate the runtime statistics of the tasks under a specific parallel strategy and workload. Then we follow the guidelines in the Megatron-LM [16] paper to prune the design space and brute-force search the optimal strategy.
Stage fusion. For inter-stage fusion, we use the simulator mentioned above to choose the best migration threshold $R_{t}$ to minimize the fused execution time under a specific workload pattern. For intra-stage fusion, we use MPI [35] to parallelize the simulated annealing with different random seeds. After finding the optimal solution, it will generate the sequences of NCCL [36] operations on each device following the schedule. The runtime will follow the instruction flow to execute the fused pipeline schedule.
System optimizations. Additionally, we adopt a series of system optimizations tailored for RLHF training. These optimizations are implemented in our own system as well as open-source RLHF frameworks like ReaLHF [13]. To separate the differences in low-level system implementations from the core techniques (inter-stage fusion in §4 and intra-stage fusion in §5), our evaluation in §7 uses RLHFuse-Base that includes these optimizations but without inter- and intra-stage fusion as an additional baseline.
@@
Task switching. The primary overhead during task switching lies in model weights reloading. For the Actor and Critic models, the latest weights after the training stage need to be redistributed to the appropriate devices according to the new parallel strategy for generation or inference. This communication pattern is well-studied [40, 14, 13] and we adopt a similar all-to-all approach to minimize cross-node communication. For the Ref and RW models whose weights remain unchanged, we keep them in CPU memory and swap them into GPU memory as needed. The swapping can overlap with the computation of previous tasks. With these optimizations, the task switching overhead is negligible compared to the total iteration time, as shown in §7.2.
@@ 7.1 End-to-End Results
7.2 Performance Analysis
To further understand the performance improvement of RLHFuse, we show its RLHF iteration breakdown and compare it with RLHFuse-Base. As shown in Figure 8, each row represents a generation setting, and each column represents a model setting X/Y. We divide one RLHF iteration into three main parts: the generation plus inference stage (Gen.+Inf.), the training stage (Train), and all other overheads (Others), such as the data transmission and weights redistribution time.
@@ 7.2 Performance Analysis
7.3 Effectiveness of Stage Fusion
In this section, we show the effectiveness of RLHFuse’s key techniques in §4 and §5.
@@ 7.3 Effectiveness of Stage Fusion
7.4 Deep Dive
In this section, we deep dive into RLHFuse by showing a fused pipeline schedule produced by RLHFuse for the 65B/33B model setting. The upper of Figure 10 shows the GPU execution timeline, with forward and backward tasks in light and dark colors, respectively. With intra-stage fusion, RLHFuse strategically fuses one 65B model of 16 PP stages (blue grids) with two 33B models that each has 8 PP stages (pink grids) to fill the pipeline bubbles of each other in reversed pipeline directions. Ultimately, the execution time of this fused pipeline schedule is the same as individually executing the 65B model using 1F1B schedule, which means we completely overlap the training of the 33B model, achieving the theoretical lower bound. Meanwhile, as shown in the bottom of Figure 10, the peak activation memory usage also achieves the lower bound of executing the two models in serial with 1F1B schedule. This complex schedule, which achieves optimal performance in both time and memory usage, is difficult to design manually, further demonstrating the effectiveness of our fused pipeline schedule generation algorithm proposed in §5.2.